OpenStudy (anonymous):

Julio has started his very own floral business. He sells centerpieces for $19.50 each. His fixed costs are $500 per month, and each centerpiece cost $8 to produce. Determine your break-even point. Suppose you wish to break even by selling only 25 centerpieces. At what price would you need to sell each centerpiece? a. Julio must sell 25 centerpieces at $35 each to break even. b. Julio must sell 25 centerpieces at $25 each to break even. c. Julio must sell 25 centerpieces at $40 each to break even. d. Julio must sell 25 centerpieces at $28 each to break even.

OpenStudy (anonymous):

also Julio has started his very own floral business. He sells centerpieces for $19.50 each. His fixed costs are $500 per month, and each centerpiece cost $8 to produce. Let n represent the number of centerpieces you sell. Write an equation to determine the cost, C. Write another equation to determine the revenue, R.

OpenStudy (anonymous):

also Several students have a really great business plan and decide to start a graphic T-shirt company. After initial expenses of $280, they will purchase each T-shirt wholesale for $3.99. They will then sell the each shirt for $10.99. How many must they sell to break even? a. 25 shirts b. 70 shirts c. 40 shirts d. 18 shirts

OpenStudy (anonymous):

also Several students decide to start a T-shirt company. After initial expenses of $280, they purchase each T-shirt wholesale for $3.99. They sell each T-shirt for $10.99. How many must they sell to break even? a. 10 shirts c. 30 shirts b. 20 shirts d. 40 shirts

OpenStudy (anonymous):

also Mary’s yard is a mess. She needs to hire someone to prune her trees and shrubs. A landscaping service she calls quotes her a price of $15 consultation fee plus $8 an hour for the actual work. Mary’s neighbor has offered to help her out. She doesn’t charge a consultation fee, but does charge $10 an hour for her work. Who do you think Mary should hire for a 3-hour job and why? a. Mary should hire the landscaping service; 3 hours of labor would cost $23, while the neighbor would charge $30. b. Mary should hire the landscaping service; 3 hours of labor would cost $30, while the neighbor would charge $39. c. Mary should hire her neighbor; 3 hours of labor would cost $39, while the landscaping service would charge $30. d. Mary should hire her neighbor; 3 hours of labor would cost $30, while the landscaping service would charge $39.

OpenStudy (anonymous):

also Suppose you invest $10,410 in equipment to manufacture a new board game. Each game costs $2.65 to manufacture and sells for $20. How many games must you make and sell before your business breaks even? a. 500 games c. 600 games b. 550 games d. 650 games

OpenStudy (anonymous):

also Erin has developed a new lotion that will help children with sensitive skin. She invests $15,225 in equipment to manufacture and package her lotion. Each bottle costs $3.25 to manufacture and sells for $12. How many bottles of lotion must she make and sell before her business breaks even? a. 998 bottles b. 1740 bottles c. 174 bottles d. 1269 bottles

OpenStudy (anonymous):

also Mary’s yard is a mess. She needs to hire someone to prune her trees and shrubs. A landscaping service she calls quotes her a price of $15 consultation fee plus $8 an hour for the actual work. Mary’s neighbor has offered to help her out. She doesn’t charge a consultation fee, but does charge $10 an hour for her work. Write two equations, one that describes the landscaping service charge and one that describes your neighbor’s charge. Let C equal the charge as a function of h, the number of hours worked. a. C = 15 + 10h C = 8h b. C = 10h + 8h C = 15 c. C = 15 + 8h C = 10h d. C = 18h C = 15

OpenStudy (anonymous):

if anyone can answer any of those i would be so greatful